OpenStudy (anonymous):

Trail Mix Recipe 1 ½ cups of peanuts 2/3 cup of M&Ms ½ cup of raisins 1/3 cup of cashews ¼ cup of sunflower seeds ¼ cup of coconut Measure all ingredients and mix in bowl. How many half cup servings will this recipe make? A) 9 servings B) 6 servings C) 7 servings D) 3.5 servings

OpenStudy (anonymous):

gotta add them all up, then divide by \(\frac{1}{2}\) which is the same as multiplying by \(2\)

OpenStudy (anonymous):

lets add smart, \(\frac{1}{3}+\frac{1}{3}=1\)

OpenStudy (anonymous):

\(1\tfrac{1}{2}+\frac{1}{2}=2\) so now we are up to \(3\)

OpenStudy (anonymous):

finally \(\frac{1}{4}+\frac{1}{4}=\frac{1}{2}\) making a total of \(3\tfrac{1}{2}\) double this to get your answer

OpenStudy (anonymous):

@satellite73 So would I multiply 3 1/2 by 2 to get my answer ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

because dividing by \(\frac{1}{2}\) is the same as multiplying by \(2\)

OpenStudy (anonymous):

@satellite73 Thank you. Also please help me with a question on @Hayleexx .... that is also my account. I have an open question I need help with.

OpenStudy (anonymous):

can you repost it? i can't find it

OpenStudy (anonymous):

@satellite73 It is the 2nd question in 'Open Questions'...it's by @Hayleexx